ABCD is a rhombus. The measure of ∠B is (x+5), and the measure of ∠C is 2x

Which statement explains why the equation 2x+x+5=180 can be used to solve for x?.


A. Opposite sides in a rhombus are parallel.
B. A square is a type of rhombus.

C. All sides in a rhombus are congruent.
D. Opposite angles in a rhombus are congruent.

Answer by mathsolverplus(88)


Given that the interior angles of any quadrilateral add up to 360,
Choice D would result in the following:
(x+5)+(x+5)+(2x)+(2x) = 360
2(x+5) + 2(2x) = 360
(x+5) + 2x = 180



If this answer has helped you and you need help on additional problems, please follow us on Instagram or message us via social media! We also help with SAT Math Questions!

***Instagram (@mathsolver.plus): https://www.instagram.com/mathsolver.plus/
Twitter (@MathSolverPlus): https://twitter.com/MathSolverPlus
Snapchat (@mathsolverplus)
Email: mathsolverplus@gmail.com